# Structural Alignment Improves Graph Test-Time Adaptation

### v1

@@ Abstract
Graph-based learning has achieved remarkable success in domains ranging from recommendation to fraud detection and particle physics by effectively capturing underlying interaction patterns. However, it often struggles to generalize when distribution shifts occur, particularly those involving changes in network connectivity or interaction patterns. Existing approaches designed to mitigate such shifts typically require retraining with full access to source data, rendering them infeasible under strict computational or privacy constraints. To address this limitation, we propose a test-time structural alignment (TSA) algorithm for Graph Test-Time Adaptation (GTTA), a novel method that aligns graph structures during inference without revisiting the source domain. Built upon a theoretically grounded treatment of graph data distribution shifts, TSA integrates three key strategies: an uncertainty-aware neighborhood weighting that accommodates structure shifts, an adaptive balancing of self-node and neighborhood-aggregated representations driven by node representations’ signal-to-noise ratio, and a decision boundary refinement that corrects remaining label and feature shifts. Extensive experiments on synthetic and real-world datasets demonstrate that TSA can consistently outperform both non-graph TTA methods and state-of-the-art GTTA baselines.
Machine Learning, ICML
Figure 1: Example of the distribution shifts of neighborhood information due to different experimental conditions in the LHC experiments. The goal of pileup mitigation is to detect LC neutral particles. The particles are modeled by kNN graphs (dashed circles in the figure) to use nearby particles for drawing inference. The model is trained on ${\mathcal{G}}^{\mathcal{S}}$ but needs to generalize to ${\mathcal{G}}^{{\mathcal{T}}_{1}}$ and ${\mathcal{G}}^{{\mathcal{T}}_{2}}$ over time. The inferred nodes within the circles are the LC neutral particles, but their neighborhood node label ratio changes in ${\mathcal{G}}^{{\mathcal{T}}_{1}}$, ${\mathcal{G}}^{{\mathcal{T}}_{2}}$. In ${\mathcal{G}}^{{\mathcal{T}}_{1}}$ the homophily ratio changes as one of the neighbor node is an LC neutral particle. Both cases represent neighborhood shift which this work aims to address. A more formal definition of these shifts is discussed in Sec. 3.
@@ 1 Introduction
Graph-based methods have become indispensable in handling structured data across a wide range of real-world applications (Duvenaud et al., 2015; Bronstein et al., 2017; Zhang et al., 2019a; Stokes et al., 2020), achieving significant success when training and testing data originate from similar distributions. However, these methods struggle to generalize to test-time data in a different domain, where variations in time, location, or experimental conditions result in distinct graph connection patterns. Although some literature on graph domain adaptation (GDA) (Wu et al., 2020; You et al., 2023; Zhu et al., 2021; Liu et al., 2023) seeks to bridge these gaps by aligning labeled source distributions with target distributions, such approaches are often infeasible in practice due to the cost of retraining and the limited availability of source data in the test time.
Figure 2: TSA utilizes neighborhood alignment and SNR-inspired adjustment to mitigate neighborhood information discrepancy. It further adjusts the decision boundary to get refined predictions $\hat{y}$. The refined soft pseudo-labels are used to estimate parameter $\bm{\gamma}$ for neighborhood alignment and to optimize $\alpha$ for combining self representations with neighborhood aggregated representations.
@@
Real-world applications may operate under constraints such as lightweight computation, limited storage, and strict privacy requirements, which make reprocessing of source data generally infeasible (Wang et al., 2021c; Wu et al., 2021). For example, in particle tracking based on graph learning at the Large Hadron Collider (LHC) (Shlomi et al., 2020; Highfield, 2008; Miao et al., 2024), the connections between particles differ as experimental conditions change over time. The tracker is expected to be capable of adapting on-the-fly (Li et al., 2023; Komiske et al., 2017). Similarly, in fraud detection for financial networks  (Clements et al., 2020; Wang et al., 2021b), privacy-sensitive training data may not be accessed during testing, while transaction patterns may change between different regions  (Wang et al., 2019; Dou et al., 2020). These scenarios indicate the need for graph test-time domain adaptation (GTTA), which enables models to adjust to new test domains without re-accessing the source domain or incurring the overhead of full retraining.
@@
Accurate node classification in graph-structured data relies heavily on effectively leveraging neighborhood information. Graph neural networks (GNNs) (Kipf & Welling, 2017; Veličković et al., 2018; Hamilton et al., 2017; Hamilton, 2020) have achieved significant success across a wide range of applications by utilizing this neighborhood context, yet they remain sensitive to distribution shifts in node neighborhoods (Gui et al., 2022; Ji et al., 2023). As shown in Fig. 1, shifts in LHC data can lead to distinct neighborhood connections among center particles in the same class (neighborhood shift). Existing methods for test-time adaptation (TTA) often focus on adjusting only the classifier’s final layer (Iwasawa & Matsuo, 2021), modifying its outputs (Boudiaf et al., 2022), or adapting normalization statistics (Wang et al., 2021a; Niu et al., 2023), primarily targeting image-based tasks with independently labeled data points. Consequently, they may not adequately address neighborhood shifts for node classification in graphs. While recent works on GTTA attempt to address the structural changes using graph augmentation (Jin et al., 2023), self-supervised learning (Mao et al., 2024). These approaches mainly rely on heuristics and homophily consistency, failing to fully address changes in neighboring node connections in principle.
@@
In this work, we propose a model-agnostic framework with test-time structural alignment, named TSA, as illustrated in Fig. 2, and supported by theoretical analysis. We begin by examining the generalization gap in GTTA for a pretrained source model. Our theoretical findings, further validated by empirical studies on synthetic datasets, indicate the importance of aligning the label distributions of neighboring nodes among the same class of nodes across the source and target domains. In the context of training-time GDA, PairAlign (Liu et al., 2024) employs a neighborhood weighting strategy to recalibrate the influence of neighboring nodes during message aggregation for a similar alignment. However, extending this approach to GTTA necessitates our more nuanced investigation, leading to the following contributions:
@@
Firstly, the weight assignment relies on knowledge of node labels, which are unavailable in the target graph under GTTA. Assigning weights based on pseudo labels may, in fact, degrade performance. To mitigate this, TSA introduces an uncertainty-aware assignment strategy that aligns only node pairs with more reliable test-time pseudo labels.
Additionally, TSA optimizes the test-time combination of self-node representations and neighborhood-aggregated representations based on their signal-to-noise ratio (SNR). For instance, when the target graph is sparse, TSA reduces reliance on neighborhood messages due to their high variance and low SNR. Conversely, as layer depth increases, it assigns greater weight to self-node representations, which become progressively denoised.
@@
Lastly, TSA integrates non-graph TTA methods to refine the decision boundary, mitigating mismatches caused by additional label and feature shifts once neighborhood shift has been addressed.
@@
We conduct extensive experiments on synthetic and four real-world datasets, including those from high-energy physics and citation networks, with multiple GNN backbones, to demonstrate the effectiveness of TSA. TSA outperforms non-graph TTA baselines by up to 12% on synthetic datasets and surpasses all non-graph baselines. Compared to existing GTTA baselines, TSA achieves an average of 10% improvement on all real-world datasets.
@@ 2.1 Notations and Problem Setup
We use upper-case letters, such as $Y$ to represent random variables, lower-case letters, such as $y$ to represent their realization. The calligraphic letters, such as ${\mathcal{Y}}$ denote the domain of random variables. We use bold capital such as ${\mathbf{Y}}$ to represent the vectorized corresponds, i.e., a collections of random variables. The probability distribution of a random variable $Y$ for a realization is expressed as $\mathbb{P}(Y=y)$.
@@
We let ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ denote an undirected and unweighted graph with the symmetric adjacency matrix ${\bm{A}}\in\mathbb{R}^{N\times N}$ and the node feature matrix ${\bm{X}}=[x_{1},\dots,x_{N}]^{T}$. GNNs utilize neighborhood information by encoding ${\bm{A}}$ and ${\bm{X}}$ into node representations $\{h_{v}^{(k)},v\in\mathcal{N}_{u}\}$. With $h_{u}^{(1)}=x_{u}$, the message passing in standard GNNs for node $v$ and each layer $k\in[L]\coloneqq\{1,\dots,L\}$ can be written as
@@
where $\mathcal{N}_{u}$ denotes the set of neighbors of node $u$, which $|\mathcal{N}_{u}|$ represent the node degree $d_{u}$. The AGG function aggregates messages from the neighbors, and the UPT function updates the node representations.
@@
Assume the model consists of a GNN feature encoder $\phi:{\mathcal{X}}\rightarrow{\mathcal{H}}$ and a classifier $g:{\mathcal{H}}\rightarrow{\mathcal{Y}}$. The model is trained on source graph ${\mathcal{G}}^{{\mathcal{S}}}=({\mathcal{V}}^{{\mathcal{S}}},{\mathcal{E}}^{{\mathcal{S}}})$ with node labels $\mathbf{y}^{\mathcal{S}}$ and the goal is to enhance model performance on test graph ${\mathcal{G}}^{{\mathcal{T}}}=({\mathcal{V}}^{{\mathcal{T}}},{\mathcal{E}}^{{\mathcal{T}}})$ with distribution shifts that will be defined in Sec. 3.1. For node classification tasks, we aim to minimize the test error $\varepsilon^{\mathcal{T}}(g\circ\phi)=\mathbb{P}^{\mathcal{T}}(g(\phi(X_{u},{\mathbf{A}}))\neq Y_{u})$ without accessing ${\mathcal{G}}^{{\mathcal{S}}}$.
@@ 2.2 Related Works
Test-time Adaptation Test-time training (Sun et al., 2020; Liu et al., 2021; Bartler et al., 2022) adapts the source model to the target domain but requires to first add a customized self-supervised losses in model pretraining. In contrast, our setup falls into the category of fully test-time adaptation (Wang et al., 2021a; Liang et al., 2024), where we do not alter the model training pipeline. Tent (Wang et al., 2021a) adapts the batch normalization parameters by minimizing the entropy, which motivated following-up studies on normalization layer adaptation  (Gong et al., 2022; Zhao et al., 2023; Lim et al., 2023; Niu et al., 2023). Some TTA works directly modify the classifier’s prediction, such as LAME (Boudiaf et al., 2022) that directly refines the model’s soft prediction by applying a regularization in the feature space, T3A (Iwasawa & Matsuo, 2021) that classifies test data based on the distance to the pseudo-prototypes derived from pseudo-labels, TAST (Jang et al., 2022) and PROGRAM (Sun et al., 2024) that extend T3A through constructing more reliable prototype graphs. However, the above methods are designed for image-based applications and cannot handle the shifts in neighborhood information of graph data.
Graph Test-time Adaptation Studies on GTTA are in two categories - node and graph classification. Graph classification problems can treat each graph as an i.i.d. input, allowing more direct extension of image-based TTA techniques to graphs (Chen et al., 2022; Wang et al., 2022). Our work focuses on node classification. GTrans (Jin et al., 2023) proposes to augment the target graph at the test time by optimizing a contrastive loss by generating positive views from DropEdge (Rong et al., 2020) and negative samples from the features of the shuffling nodes (Veličković et al., 2019). GraphPatcher (Ju et al., 2024) learns to generate virtual neighbors to improve low-degree nodes classification. SOGA (Mao et al., 2024) designs a self-supervised loss that relies on mutual information maximization and homophily assumption. These works are mostly built upon heuristics and may not address structure shifts in principle.
@@ 3 Test Error Analysis
In this section, we characterize the generalization error between source and target graphs and explicitly attribute it to three different kinds of shifts: label shift, feature shift, as well as neighborhood shift. Motivated by our theoretical analysis on the generalization error, we then propose TSA algorithm to minimize the across-graph generalization error.
(a) Reliability of $\gamma$ (b) Source (c) Nbr. Align (d) SNR $\alpha$

### v2

@@ Abstract
Graph-based learning excels at capturing interaction patterns in diverse domains like recommendation, fraud detection, and particle physics. However, its performance often degrades under distribution shifts, especially those altering network connectivity. Current methods to address these shifts typically require retraining with the source dataset, which is often infeasible due to computational or privacy limitations. We introduce Test-Time Structural Alignment (TSA), a novel algorithm for Graph Test-Time Adaptation (GTTA) that aligns graph structures during inference without accessing the source data. Grounded in a theoretical understanding of graph data distribution shifts, TSA employs three synergistic strategies: uncertainty-aware neighborhood weighting to accommodate neighbor label distribution shifts, adaptive balancing of self-node and aggregated neighborhood representations based on their signal-to-noise ratio, and decision boundary refinement to correct residual label and feature shifts. Extensive experiments on synthetic and real-world datasets demonstrate TSA’s consistent outperformance of both non-graph TTA methods and state-of-the-art GTTA baselines.
@@ 1 Introduction
Graph-based methods have become indispensable in handling structured data across a wide range of real-world applications [1, 2, 3, 4], achieving significant success when training and testing data originate from similar distributions. However, these methods struggle to generalize to test-time data in a different domain, where variations in time, location, or experimental conditions result in distinct graph connection patterns. Although some literature on graph domain adaptation (GDA) [5, 6, 7, 8] seeks to bridge these gaps by aligning labeled source distributions with target distributions, such approaches are often infeasible in practice due to the cost of retraining and the limited availability of source data at the test time.
@@
Real-world applications may operate under constraints such as lightweight computation, limited storage, and strict privacy requirements, which make reprocessing of source data generally infeasible [9, 10]. For example, Fig. 1 illustrates pileup mitigation based on graph learning at the Large Hadron Collider (LHC) [11, 12, 13], where the connections between particles differ as experimental conditions change. To address this, the detector is expected to be capable of adapting on-the-fly [14, 15], potentially on embedded hardware. Similarly, in fraud detection for financial networks [16, 17], privacy-sensitive training data may not be accessed during testing, while transaction patterns may change between different regions [18, 19]. These scenarios indicate the need for graph test-time domain adaptation (GTTA), which enables models to adjust to new test domains without re-accessing the massive source dataset during adaptation or incurring the overhead of full retraining.
@@
Accurate node classification in graph-structured data relies heavily on effectively leveraging neighborhood information. Graph neural networks (GNNs) [20, 21, 22, 23] have achieved significant success across a wide range of applications by utilizing this neighborhood context, yet they remain sensitive to distribution shifts in node neighborhoods [24, 25]. In the above LHC example illustrated by Fig. 1, distribution shifts lead to distinct neighborhood connections among center particles in the same class, causing a neighborhood shift. Existing methods for test-time adaptation (TTA) often focus on adjusting only the classifier’s final layer [26], modifying its outputs [27], or adapting normalization statistics [28, 29], primarily targeting image-based tasks with independently labeled data points. Consequently, they may not adequately address neighborhood shifts for node classification in graphs. While recent works on GTTA attempt to address the structural changes using graph augmentation [30], self-supervised learning [31, 32], these approaches mainly rely on heuristics and homophily consistency, failing to fully address changes in neighboring node connections in principle.
@@
In this work, we propose a model-agnostic framework with test-time structural alignment, named TSA, as illustrated in Fig. 2, and supported by theoretical analysis. We begin by examining the generalization gap in GTTA for a pretrained source model. Our theoretical findings, further validated by empirical studies, indicate the importance of aligning the neighborhood distributions across the source and target domains. Our contributions are summarized as follows:
@@
Firstly, TSA employs a neighborhood weighting strategy in GTTA to recalibrate the influence of neighboring nodes during message aggregation, thereby enabling structural alignment under the GTTA setting. However, the weight assignment relies on knowledge of node labels, which are unavailable in the target graph. Assigning weights based on pseudo labels may, in fact, degrade performance. To mitigate this, TSA introduces an uncertainty-aware assignment strategy that aligns only node pairs with more reliable test-time pseudo labels.
Additionally, TSA optimizes the test-time combination of self-node representations and neighborhood-aggregated representations based on their signal-to-noise ratio (SNR). Our experiments show that neighborhood-aggregated representations tend to have higher SNR and are better denoised when the neighbor set size and the GNN layer depth increase. Under both conditions, TSA adjusts the combination by assigning greater weight to neighborhood-aggregated representations than to self-node representations to enhance model performance.
@@
Lastly, TSA integrates non-graph TTA methods to refine the decision boundary, mitigating mismatches caused by additional label and feature shifts once neighborhood shifts have been addressed.
@@
We conduct extensive experiments on synthetic and four real-world datasets, including those from high-energy physics and citation networks, with multiple GNN backbones, to demonstrate the effectiveness of TSA. TSA outperforms non-graph TTA baselines by up to 21.25% on synthetic datasets and surpasses all non-graph baselines. Compared to existing GTTA baselines, TSA achieves an average improvement of 10% on all real-world datasets.
Figure 1: Example of the distribution shifts of neighborhood information due to different experimental conditions. Pileup mitigation aims to detect LC neutral particles using kNN graphs (dashed circles) that leverage nearby particles for inference. The model is trained on ${\mathcal{G}}^{\mathcal{S}}$ but needs to generalize to ${\mathcal{G}}^{{\mathcal{T}}_{1}}$ and ${\mathcal{G}}^{{\mathcal{T}}_{2}}$. The inferred nodes within the circles are the LC neutral particles, but their neighborhood node label ratios change in ${\mathcal{G}}^{{\mathcal{T}}_{1}}$ and ${\mathcal{G}}^{{\mathcal{T}}_{2}}$. Both cases represent neighborhood shift, which this work aims to address. A more formal definition of these shifts is discussed in Sec. 3.
Figure 2: TSA utilizes neighborhood alignment to address neighborhood shift and SNR-inspired adjustment to mitigate SNR shift. It further adjusts the decision boundary to get refined predictions $\hat{y}$. The refined soft pseudo-labels are used to estimate the parameter $\bm{\gamma}$ for neighborhood alignment and to optimize $\alpha$ for combining self representations with neighborhood-aggregated representations.
@@ 2.1 Notations and Problem Setup
We use upper-case letters, such as $Y$ to represent random variables, and lower-case letters, such as $y$ to represent their realizations. Calligraphic letters, such as ${\mathcal{Y}}$ denote the domain of random variables. We use bold capital letters such as ${\mathbf{Y}}$ to represent the vectorized counterparts, i.e., collections of random variables. The probability of a random variable $Y$ for a realization is expressed as $\mathbb{P}(Y=y)$.
@@
We let ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ denote an undirected and unweighted graph with the symmetric adjacency matrix ${\bm{A}}\in\mathbb{R}^{N\times N}$ and the node feature matrix ${\bm{X}}=[x_{1},\dots,x_{N}]^{T}$. GNNs utilize neighborhood information by encoding ${\bm{A}}$ and ${\bm{X}}$ into node representations $\{h_{v}^{(k)},v\in\mathcal{N}_{u}\}$. With $h_{u}^{(1)}=x_{u}$, the message passing in standard GNNs for node $u$ and each layer $k\in[L]\coloneqq\{1,\dots,L\}$ can be written as
@@
where $\mathcal{N}_{u}$ denotes the set of neighbors of node $u$, and $|\mathcal{N}_{u}|$ represents the node degree $d_{u}$. The AGG function aggregates messages from the neighbors, and the UPT function updates the node representations.
Test-Time Adaptation (TTA).
Assume the model consists of a feature encoder $\phi:{\mathcal{X}}\rightarrow{\mathcal{H}}$ and a classifier $g:{\mathcal{H}}\rightarrow{\mathcal{Y}}$ and each domain ${\mathcal{U}}\in\{{\mathcal{S}},{\mathcal{T}}\}$ consists of a joint feature and label distribution $\mathbb{P}^{\mathcal{U}}(X,Y)$. In TTA, the model is first trained on a labeled source dataset $\{(x_{i},y_{i})\}_{i=1}^{N}$, IID sampled from the source domain. In the target dataset $\{x_{i}\}_{i=1}^{M}$, the objective is to minimize the test error $\varepsilon^{\mathcal{T}}(g\circ\phi)=\mathbb{P}^{\mathcal{T}}(g(\phi(X))\neq Y)$ without accessing the source dataset during adaptation.
@@
Assume the model consists of a GNN feature encoder $\phi:{\mathcal{X}}\rightarrow{\mathcal{H}}$ and a classifier $g:{\mathcal{H}}\rightarrow{\mathcal{Y}}$. The model is trained on source graph ${\mathcal{G}}^{{\mathcal{S}}}=({\mathcal{V}}^{{\mathcal{S}}},{\mathcal{E}}^{{\mathcal{S}}})$ with node labels $\mathbf{y}^{\mathcal{S}}$ and the goal is to enhance model performance on a test graph ${\mathcal{G}}^{{\mathcal{T}}}=({\mathcal{V}}^{{\mathcal{T}}},{\mathcal{E}}^{{\mathcal{T}}})$ with distribution shifts that will be defined in Sec. 3.1. For node classification tasks, we aim to minimize the test error $\varepsilon^{\mathcal{T}}(g\circ\phi)=\mathbb{P}^{\mathcal{T}}(g(\phi(X_{u},{\mathbf{A}}))\neq Y_{u})$ without accessing ${\mathcal{G}}^{{\mathcal{S}}}$ during adaptation.
@@ 2.2 Related Works
Test-time Adaptation.
Test-time training (TTT) approaches necessitate modifications to both the training pipeline and network architecture to accommodate customized self-supervised tasks, such as rotation prediction [33], contrastive learning [34, 35], and clustering [36]. These requirements, however, can limit their broader applicability. In contrast, our proposed setup aligns with the test-time adaptation (TTA) category, which does not alter the original training pipeline [37]. One prominent TTA method, Tent [28], adapts batch normalization parameters by minimizing entropy. This approach has inspired subsequent research focusing on normalization layer adaptation [38, 39, 40, 29]. Other TTA strategies directly modify the classifier’s predictions. For instance, LAME [27] refines the model’s soft predictions by applying regularization in the feature space. T3A [26] classifies test data based on its distance to pseudo-prototypes derived from pseudo-labels. Building upon T3A, TAST [41] and PROGRAM [42] aim to construct more reliable prototype graphs. However, the effectiveness of these methods can be compromised when significant domain shifts occur [43, 44]. To address larger domain shifts, some methods leverage stored, lightweight source statistics, such as label distributions, to perform test-time alignment [43, 45, 46, 47, 48, 49]. ActMAD [43], for example, aligns target feature statistics with stored source statistics across multiple layers. More recently, FOA [45] focuses on aligning the statistics of classification tokens in vision transformers. Note that all the aforementioned methods are primarily designed for image-based applications and may not adequately address shifts in the neighborhood information characteristic of graph data.
Graph Test-time Adaptation.
Studies on GTTA are in two categories - node and graph classification. Graph classification problems can treat each graph as an i.i.d. input, allowing more direct extension of image-based TTA techniques to graphs [50, 51]. Our work focuses on node classification [30, 31, 32, 52]. GTrans [30] proposes to augment the target graph at test time by optimizing a contrastive loss by generating positive views from DropEdge [53] and negative samples from the features of the shuffled nodes [54]. HomoTTT [31] combines self-supervised learning with homophily-based DropEdge and prediction selection. SOGA [32] designs a self-supervised loss that relies on mutual information maximization and the homophily assumption. These works are mostly built upon heuristics and may not address structure shifts in principle.
@@ 3 Test Error Analysis
In this section, we characterize the generalization error between the source and target graphs and explicitly attribute it to different types of shifts in graph data, which further motivate our TSA algorithm.
3.1 Distribution Shifts on Graphs
Distribution shifts on graphs have been formally studied in previous GDA works [5, 6, 7, 55, 56, 57]. Following their definitions, we categorize shifts in graphs into two types: feature shift and structure shift. For simplicity, our analysis is based on a data generation process: ${\mathbf{X}}\leftarrow{\mathbf{Y}}\rightarrow{\mathbf{A}}$, where the graph structure and node features are both conditioned on node labels.
Definition 3.1.
(Feature Shift). Assume node features $x_{u},u\in{\mathcal{V}}$ are i.i.d sampled given labels $y_{u}$, then we have $\mathbb{P}({\mathbf{X}}|{\mathbf{Y}})=\prod_{u\in{\mathcal{V}}}\mathbb{P}(X_{u}|Y_{u})$. We then define feature shift as $\mathbb{P}^{\mathcal{S}}(X_{u}|Y_{u})\neq\mathbb{P}^{\mathcal{T}}(X_{u}|Y_{u})$.
Definition 3.2.
(Structure Shift). As the graph structure involves the connecting pattern between nodes including their labels, we consider the joint distribution of the adjacency matrix and labels $\mathbb{P}({\mathbf{A}},{\mathbf{Y}})$, where structure shift, denoted by $\mathbb{P}^{\mathcal{S}}({\mathbf{A}},{\mathbf{Y}})\neq\mathbb{P}^{\mathcal{T}}({\mathbf{A}},{\mathbf{Y}})$, can be decomposed into conditional structure shift (CSS) $\mathbb{P}^{\mathcal{S}}({\mathbf{A}}|{\mathbf{Y}})\neq\mathbb{P}^{\mathcal{T}}({\mathbf{A}}|{\mathbf{Y}})$ and label shift (LS) $\mathbb{P}^{\mathcal{S}}({\mathbf{Y}})\neq\mathbb{P}^{\mathcal{T}}({\mathbf{Y}})$. Here, CSS directly reflects the shift in neighborhood information (neighborhood shift) in Fig. 1.
3.2 Theoretical Analysis
Let $\varepsilon^{\mathcal{S}}(g\circ\phi)$ denote the error of a pretrained GNN on the source domain and $\varepsilon^{\mathcal{T}}(g\circ\phi)$ the error of the model when applied to a test graph in the target domain. Inspired by Tachet des Combes et al. [58], we provide an upper bound on the error gap between the source and target domains, showing how a pretrained GNN (e.g., ERM) can be influenced. We denote the balanced error rate of a pretrained node predictor $\hat{Y}_{u}$ on the source domain as $\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})\coloneq\max_{j\in\mathcal{Y}}\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}\neq Y_{u}|Y_{u}=j).$
Theorem 3.3 (Error Decomposition Theorem (Informal)).
Suppose ${\mathcal{G}}^{{\mathcal{S}}}$ and ${\mathcal{G}}^{{\mathcal{T}}}$, we can decouple both graphs into independent ego-networks (center nodes and 1-hop neighbors). For any classifier $g$ with a GNN encoder $\phi$ in node classification tasks, we have the following upper bound on the error gap between source and target under feature shift and structure shift:
$\leq\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})\cdot\{\underbrace{(2\text{TV}(\mathbb{P}^{\mathcal{S}}(Y_{u}),\mathbb{P}^{\mathcal{T}}(Y_{u}))}_{\text{Label Shift}}$ $+\underbrace{\mathbb{E}_{Y_{u}}^{\mathcal{T}}[\max_{k\in{\mathcal{Y}}}\lvert 1-\frac{\mathbb{P}^{\mathcal{T}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}{\mathbb{P}^{\mathcal{S}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}\rvert]}_{\textbf{Conditional Structure Shift}}\}+\underbrace{\Delta_{CE}}_{\text{Feature Shift}}$
Our bound aligns with the findings of Liu et al. [59], which highlight the impact of conditional structure shift, label shift, and feature shift on generalization to the target graph. We extend this understanding by deriving an explicit error bound.
4 Test-Time Structural Alignment (TSA)
(a) Reliability of $\gamma$ (b) Source (c) Target with Stru. Shift (d) Nbr. Align